### v1

@@ 8.3 Topological Statistics
Persistence images are stable vector representations of persistence diagrams (Adams et al., 2017), and they were used by Thorne et al. (2022) as summary statistics to infer two parameters in the AC model. To obtain persistence images, persistence pairs $(b,d)$ are first transformed to the modified pair $(b,d-b)$. A persistence surface is then the function $\Xi:\mathbb{R}^{2}\to\mathbb{R}$ (24), where the sum is taken over all modified pairs.
@@
Here, $w$ is a weighting function on the modified persistence pairs, and we use the standard choice $w(b,d-b)=d-b/\max{(d-b)}$, which divides the persistence by the maximum persistence across all persistence pairs. The function $\Phi_{(b,d-b)}$ must be a differentiable probability distribution with mean $(b,d-b)$, and we take the standard Gaussian distribution with variance $0.1$. A persistence image is then a collection of integrals (25) of $\Xi$ over discretised regions $\mathcal{R}\subset\mathbb{R}^{2}$:
@@
The regions $\mathcal{R}$ must be chosen so that the integrals reflect the distribution of persistence pairs used to construct $\Xi$. The persistence pairs we compute from angiogenesis datasets with sweeping-plane filtrations have birth and death values between $k=0$ and $k=200$. We therefore take regions $R_{i,j}=[20i,20(i+1)]\times[20j,20(j+1)]$ for $i,j=0,1,\dots,9$ to generate a total of $100$ integrals which constitute a persistence image. For each angiogenesis dataset we compute eight such persistence images–one for each of the four types of extended persistence pair described in Section 8.2 computed using each sweeping-plane filtration (horizontal and vertical), giving a total of $800$ topological statistics.
Persistence statistics are simple summaries of EPDs that have been found to perform well on well-known classification tasks (Ali et al., 2023). To obtain persistence statistics, we consider the births $b$, deaths $d$, persistences $d-b$ and midpoints $(b+d)/2$ of persistence pairs $(b,d)$ in an EPD, and compute the mean, standard deviation, median, interquartile range, full range, and the 10th, 25th, 75th and 90th percentiles of each of these descriptors. We therefore compute a total of $36$ persistence statistics for each of the four types of persistence pairs described in Section 8.2 for each sweeping-plane filtration (horizontal and vertical), for a total of $288$.
@@ 8.4 Example Computation
We compute spatially-averaged and topological statistics from angiogenesis datasets and concatenate these into a long-list of summary statistics. Each angiogenesis dataset yields $20$ spatially-averaged statistics and $800+288$ topological statistics, which we combine into a vector of length $1108$. An example of the computation of spatially-averaged statistics and EPDs for a simple angiogenesis dataset is given in Figure 5.
@@
Figure 5: An example model simulation and computation of spatially-averaged statistics and EPDs. A) Data pre-processing. We display the final time-step of a simple simulation in which two ECs migrate from the bottom of the simulation domain, branch, anastomose and form a loop in their path upward towards the tumour. We convert the final simulation time-step into an angiogenesis dataset by overlaying a grid of $200\times 200$ pixels onto the simulation domain and colouring pixels which correspond to the location of simulated ECs. Simplicial complexes $\Sigma_{k}$ contain a vertex for each pixel location containing an EC whose $x$ co-ordinate (or $y$ coordinate) is less than or equal to $k/200$–we plot $\Sigma_{K}$ where $K=200$. B) Spatially-averaged statistics. Spatially-averaged statistics consider the $(x,y)$ co-ordinates of pixel locations which contain ECs. We compute the mean, standard deviation, minimum, maximum, interquartile range, range, 10th percentile, 25th percentile, 75th percentile and 90th percentile in each co-ordinate direction. C) EPDs. We compute EPDs to quantify the size and location of the loops (), the branches ($\boldsymbol{\times}$ and $\boldsymbol{+}$), and components ( ) at the final time-step, where the size and location are measured in either the horizontal ($x$) or vertical ($y$) direction. Computing EPH via the sequences (21) and (23) using the horizontal and vertical filtration gives two EPDs containing persistence pairs $(b,d)$ which correspond to the birth and death of topological features within the concatenated sequence. The values $b$ and $d$ correspond to the location (in co-ordinate directions) of each feature and the persistence $d-b$ gives the size. As the final step in computing topological statistics, we vectorise EPDs into persistence images and persistence statistics according to Section 8.3. The left panel of A was created in https://BioRender.com
@@ 9 ABC and Random Forest Methodology
We use Random Forests (Breiman, 2001) to choose informative summary statistics, which we then use to infer spatial parameters and select between spatial models using ABC. In this section, we give further details on each step described in Section 3 of the main text.
@@ 9.1 Step 1: Identify Informative Summary Statistics For Each Model
ABC requires a distance function $\nu:\mathcal{D}\times\mathcal{D}\to\mathbb{R}$ to measure the discrepancy between observed data and simulated data. In Section 8 we showed how to compute spatially-averaged and topological statistics from angiogenesis datasets. Combining all spatially-averaged and topological statistics into a single vector gives a total of $n_{\text{f}}=1108$ features which may be computed from an angiogenesis dataset. We now identify an informative subset of these summary statistics which we will use to construct the distance function $\nu$.
We train one regression Random Forest per parameter per angiogenesis model to learn the relationship between parameter values and summary statistics. By training one Random Forest per parameter, we aim to identify the summary statistics that capture the effect of varying each individual parameter’s value on simulated data. To generate training data, we draw $n=10,000$ parameter values $\theta_{i}$ uniformly from the parameter ranges given in Table 1. We generate an angiogenesis dataset corresponding to each parameter value by simulating each model up to time $t_{\text{final}}=4$ and consider the angiogenesis dataset at its final time-step as described in Section 7. We compute spatially-averaged and topological statistics to form a vector $X_{i}$ of summary statistics corresponding to the parameter value $\theta_{i}$, which together form training data $\mathcal{X}$ and $\mathcal{Y}$.
@@
$Q(\mathcal{N})$ is a measure of the impurity of samples allocated to node $\mathcal{N}$ and |$\absolutevalue{\mathcal{N}}$ | is the number of samples allocated to the node. As is common in regression Random Forests, we use the L2 impurity given by Equation (27) to decide splitting rules. The L2 purity measures the variance in parameter values from the mean $\bar{\theta}_{\mathcal{N}}$ among pairs allocated to the same daughter node.
@@
To choose the number $n_{\text{tree}}$ of decision trees to use in each Random Forest, we follow the advice in Raynal et al. (2018) and compute the out-of-bag mean square error of Random Forests constructed with different numbers of trees. We note in Figure 6 that the out-of-bag error decreases when adding additional trees, but that this improvement is small when using more than $100$ trees in the Random Forest for each parameter in each model. We therefore use $n_{\text{tree}}=100$.
To select a subset of the spatially-averaged and topological statistics for use in subsequent ABC algorithms, we consider the importance of each feature in the Random Forest for each parameter. The importance of a co-variate $j$ in a trained Random Forest is defined as the mean decrease in impurity $Q(\mathcal{N})-(Q(\mathcal{N}_{L})+Q(\mathcal{N}_{R}))$ achieved by all internal nodes $\mathcal{N}$ which use $j$ in their splitting rule. Features with high importance are therefore those which are most effective in partitioning the training data, meaning they capture the effect of model parameters on model simulations.
@@ 9.2 Step 2: Fit Each Model To The Observed Data
Using observed data $\mathcal{D}^{*}$, we approximate the parameter posterior $p(\Theta|\mathcal{D}^{*})$ for each angiogenesis model. To construct the distance function $\nu$ required by ABC algorithms to measure the discrepancy between simulated and observed data, we follow Section 9.1 and choose a subset of $100$ summary statistics for each model from among the spatially-averaged and topological statistics. It is possible that the important summary statistics may be poorly scaled. For example, connected components are generally larger than loops in the angiogenesis datasets we consider. Small differences in the size and location of loops in angiogenesis datasets may be stronger predictors of parameter values than the size and locations of connected components, yet this may not be reflected in the distance function $\nu$ if the corresponding topological statistics are left unscaled. We therefore construct a scaling function which divides each summary statistic by the maximum absolute value of that summary statistic among the training data for each model. The ABC distance function is then $\nu(\mathcal{D}^{*},\mathcal{D}_{i})=\norm{x^{*}-x_{i}}_{2}$ where $\mathcal{D}^{*}$ is observed data, $\mathcal{D}_{i}$ is simulated data, and $x^{*}$ and $x_{i}$ are the $100$ summary statistics with the scaling applied. When $\mathcal{D}^{*}$ comprises several instances of observed data, we average the distance $\nu$ using summary statistics computed from each.
@@
Using this distance function, we use the ABC-SMC algorithm of Del Moral et al. (2012) to approximate the parameter posterior $p(\Theta|\mathcal{D}^{*})$. ABC-SMC outputs a series of $n_{\text{pop}}$ intermediate distributions, which correspond to a decreasing sequence of tolerances $\epsilon_{0}>\dots>\epsilon_{n_{\text{pop}}}>0$. Beginning with a population of $N_{\text{pop}}$ parameters sampled from a prior distribution, the initial tolerance $\epsilon_{0}$ is chosen so that a predetermined fraction $\alpha\in(0,1)$ of parameters are accepted. The Effective Sample Size (ESS) of each population is a measure of the independence of parameters within a current population. Del Moral et al. (2012) computes the ESS of each subsequent population and chooses further tolerances such the ESS of each population decreases by $\alpha$. We use $\alpha=0.8$ and generate $n_{\text{pop}}=50$ populations, each containing $N_{\text{pop}}=1,000$ parameters, and take the final population as the approximate ABC-SMC posterior $p(\Theta|\mathcal{D}^{*})$.
@@ 9.3 Step 3: Approximate The Model Posterior
We approximate the model posterior $p(m|\mathcal{D}^{*})$ for observed data $\mathcal{D}^{*}$ by training two more Random Forests, following the method of Pudlo et al. (2015). Using the training data $(X_{i},m_{i})$ obtained from the training data used in Step 1 by replacing the parameter $\theta_{i}$ with the model index $m_{i}\in\{\text{AC},\text{SL},\text{PS}\}$, we train a classification Random Forest to learn the relationship between summary statistics $X_{i}$ and model index $m_{i}$.
Step 1 chooses a subset of the spatially-averaged and topological statistics which are important for each for each model. Since the importance of each summary statistic to a model is determined by that model’s training data, it is possible that different summary statistics are important for different models. Using summary statistics which are only informative for a single model to approximate the model posterior may bias predictions in favour of that model. We therefore learn the relationship between summary statistics and model index using only those summary statistics which are important for all angiogenesis models.
@@
The trained classification Random Forest provides a prediction $RF(X^{*})$ of the model $m^{*}$ which generated observed data $\mathcal{D}^{*}$ (which we aggregate when $\mathcal{D}^{*}$ contains multiple instances of observed data). The Random Forest also provides an out-of-bag error for each pair $(X_{i},m_{i})$, which is the proportion of those decision trees which did not use this pair in their training bootstrap which predict the incorrect model index $m_{i}$. Following Pudlo et al. (2015), we train another regression Random Forest using the same training data to learn the relationship between summary statistics $X_{i}$ and this mis-classification error rate $p(RF(X_{i})\neq m_{i})$ of the first Random Forest. Predicting the out-of-bag error $p(RF(X^{*})\neq m^{*})$ for the observed data $\mathcal{D}^{*}$ gives an estimate $1-p(RF(X^{*})\neq m^{*})$ for the model posterior $p(m=m^{*}|\mathcal{D}^{*})$.
Figure 7: Left axis: for each parameter in each model, we rank summary statistics by their importance and observe how importance decreases. Right axis: we compare how many summary statistics appear among the top $n_{\text{s}}$ most important features for all models as $n_{\text{s}}$ is increased and plot the expected number of common summary statistics if they were chosen randomly. Choosing $n_{\text{s}}=100$ includes approximately the $25$ most important summary statistics for each parameter. Selecting more summary statistics than this would include those which are of little use in capturing the effect of parameter values on simulated data. Choosing $n_{\text{s}}=100$ also ensures that $\tilde{n}_{\text{s}}=30$ summary statistics appear in the top $n_{\text{s}}$ most important summary statistics for all models. If $100$ summary statistics were chosen uniformly randomly for each model, the expected number that would be common to all three models is approximately $1$.

### v2

@@ 8.3 Topological Summary Statistics
Persistence Images are stable vector representations of persistence diagrams (persistenceimage), and they were used by tabc as summary statistics to infer two parameters in the AC model. To obtain Persistence Images, persistence pairs $(b,d)$ are first transformed to the modified pair (b,db)(b,d-b). A persistence surface is then the function $\Xi:\mathbb{R}^{2}\to\mathbb{R}$ (22), where the sum is taken over all modified pairs.
@@
Here, $w$ is a weighting function on the modified persistence pairs, and we use the standard choice $w(b,d-b)=d-b/\max{(d-b)}$, which divides the persistence by the maximum persistence across all persistence pairs. The function $\Phi_{(b,d-b)}$ must be a differentiable probability distribution with mean (b,db)(b,d-b), and we take the standard Gaussian distribution with variance $0.1$. A Persistence Image is then a collection of integrals (23) of $\Xi$ over discretised regions $\mathcal{R}\subset\mathbb{R}^{2}$:
@@
The regions $\mathcal{R}$ must be chosen so that the integrals reflect the distribution of persistence pairs used to construct $\Xi$. The persistence pairs we compute from angiogenesis datasets with sweeping-plane filtrations have birth and death values between $k=0$ and $k=200$. We therefore take regions $R_{i,j}=[20i,20(i+1)]\times[20j,20(j+1)]$ for $i,j=0,1,\dots,9$ to generate a total of $100$ integrals which constitute a Persistence Image. For each sweeping-plane filtration (horizontal and vertical), we compute four persistence images–one for each of the four types of extended persistence pair described in section 8.2 . Therefore, Persistence Images give a total of $800$ topological summary statistics for each angiogenesis dataset.
Persistence statistics are a way to summarise a collection of persistence pairs, which have been found to perform well on well-known classification tasks (vectorisationSurvey). To obtain persistence statistics, one considers the births $b$, deaths $d$, persistences dbd-b and midpoints $(b+d)/2$ of persistence pairs $(b,d)$, and computes the mean, standard deviation, median, interquartile range, full range, and the 10th, 25th, 75th and 90th percentiles of each of these quantities. Within each EPD, we consider each type of extended persistence pair (ordinary, relative, extended+, and extended-) separately, since each quantifies a different spatial feature within angiogenesis data (see Figure 6 for examples). We therefore compute a total of $36$ persistence statistics for each of the four types of persistence pairs described in section 8.2 for each sweeping-plane filtration (horizontal and vertical) Therefore, persistence statistics give a total of $36\times 4\times 2=288$ topological summary statistics for each angiogenesis dataset.
@@ 8.4 Example Computation
We compute spatially-averaged and topological summary statistics from angiogenesis datasets and concatenate these into a long-list of summary statistics. Each angiogenesis dataset yields $20$ spatially-averaged summary statistics and $800+288$ topological summary statistics, which we combine into a vector of length $1108$. An example of the computation of spatially-averaged summary statistics and EPDs for a simple angiogenesis dataset is given in Figure 6.
8.5 Distribution of Important Summary Statistics
In Figure 1 of the main text, we report the number of each type of summary statistics which were selected as informative for each model by step 1 in section 4 of the main text. Figure 7 shows the breakdown of the top 100 summary statistics for each parameter in each model by type (spatially-averaged or topological), direction (vertical or horizontal), as well as vectorisation (Persistence Images or persistence statistics). We found that a mixture of Persistence Images and persistence statistics are informative for inferring parameter values.
No spatially-averaged summary statistics computed in the horizontal ($x$) direction are selected by step 1 of our pipeline for any parameter in any model. The $x$ co-ordinates of EC locations are clustered around the four initialisation points on the horizontal axis in simulations. Some parameter values cause ECs to remain tightly clustered around these four points, and others cause them to spread out. However, horizontal spatially-averaged statistics are too coarse to distinguish these scenarios, which may explain why they are not useful in parameter inference.
@@ 8.5 Distribution of Important Summary Statistics
Figure 6: An example model simulation and computation of spatially-averaged summary statistics, persistence diagrams (PDs) and extended persistence diagrams (EPDs). A) Data pre-processing. We display the final time-step of a simple simulation in which two ECs migrate from the bottom of the simulation domain, branch, anastomose and form a loop in their path upward towards the tumour. We convert the final simulation time-step into an angiogenesis dataset by overlaying a grid of $200\times 200$ pixels onto the simulation domain and colouring pixels which correspond to the location of simulated ECs. Simplicial complexes $\Sigma_{k}$ contain a vertex for each pixel location containing an EC whose $x$ co-ordinate (or $y$ coordinate) is less than or equal to k/200k/200–we plot $\Sigma_{K}$ where $K=200$. B) Spatially-averaged summary statistics. We compute 10 simple statistics on the $x$ and $y$ co-ordinates of pixel locations which contain ECs. C)-D) PDs and EPDs. We compute persistent homology (PH) and extended persistent homology (EPH) via the sequences (19) and (21) using the horizontal and vertical filtration illustrating the additional information offered by EPH. As the final step in computing topological summary statistics, we vectorise EPDs into Persistence Images and persistence statistics according to section 8.3. The left panel of A was created in https://BioRender.com
Figure 7: Distribution of important summary statistics. Defining importance as in section 3.1 as the mean decrease in L2 impurity (25) induced by each summary statistic in each parameter’s Random Forest, we report the types of the top $100$ summary statistics. In Figure 1 of the main text, we disaggregate the important summary statistics by type (spatially-averaged or topological) and direction (horizontal (x) and vertical (y)). Here, we further delineate and show how many of each summary statistic comes from each topological vectorisation–Persistence Images and persistence statistics. No spatially-averaged summary statistics computed in the $x$ direction appear among the top $100$ for any parameter. vectorisationSurvey found that persistence statistics perform better than other vectorisation methods (including Persistence Images) on common classification tasks. We find that a mixture of Persistence Images and persistence statistics are useful for the inference of parameters in each model.
@@ 9 ABC and Random Forest Methodology
We use Random Forests (orginalRF) to choose informative summary statistics, which we then use to infer spatial parameters and select between spatial models using ABC. In this section, we give further details on each step described in section 3 of the main text.
@@ 9.1 Step 1: Identify Informative Summary Statistics For Each Model
ABC requires a distance function $\nu:\mathcal{D}\times\mathcal{D}\to\mathbb{R}$ to measure the discrepancy between observed data and simulated data. In section 8 we showed how to compute spatially-averaged and topological summary statistics from angiogenesis datasets. Combining all spatially-averaged and topological summary statistics into a single vector gives a total of $n_{\text{f}}=1108$ features which may be computed from an angiogenesis dataset. We now identify an informative subset of these summary statistics which we will use to construct the distance function $\nu$.
We train one regression Random Forest per parameter per angiogenesis model to learn the relationship between parameter values and summary statistics. By training one Random Forest per parameter, we aim to identify the summary statistics that capture the effect of varying each individual parameter’s value on simulated data. To generate training data, we draw $n=10,000$ parameter values $\theta_{i}$ uniformly from the parameter ranges given in Table 1. We generate an angiogenesis dataset corresponding to each parameter value by simulating each model up to time $t_{\text{final}}=4$ and consider the angiogenesis dataset at its final time-step as described in section 7. We compute spatially-averaged and topological summary statistics to form a vector $X_{i}$ of summary statistics corresponding to the parameter value $\theta_{i}$, which together form training data $\mathcal{X}$ and $\mathcal{Y}$.
@@
$Q(\mathcal{N})$ is a measure of the impurity of samples allocated to node $\mathcal{N}$ and $|\absolutevalue{\mathcal{N}}$ is the number of samples allocated to the node. As is common in regression Random Forests, we use the L2 impurity given by Equation (25) to decide splitting rules. The L2 purity measures the variance in parameter values from the mean $\bar{\theta}_{\mathcal{N}}$ among pairs allocated to the same daughter node.
@@
To choose the number $n_{\text{tree}}$ of decision trees to use in each Random Forest, we follow the advice in ABCRF_paraminf and compute the out-of-bag mean square error of Random Forests constructed with different numbers of trees. We note in Figure 8 that the out-of-bag error decreases when adding additional trees, but that this improvement is small when using more than $100$ trees in the Random Forest for each parameter in each model. We therefore use $n_{\text{tree}}=100$.
To select a subset of the spatially-averaged and topological summary statistics for use in subsequent ABC algorithms, we consider the importance of each feature in the Random Forest for each parameter. The importance of a co-variate $j$ in a trained Random Forest is defined as the mean decrease in impurity $Q(\mathcal{N})-(Q(\mathcal{N}_{L})+Q(\mathcal{N}_{R}))$ achieved by all internal nodes $\mathcal{N}$ which use $j$ in their splitting rule. Features with high importance are therefore those which are most effective in partitioning the training data, meaning they capture the effect of model parameters on model simulations.
@@ 9.2 Step 2: Fit Each Model To The Observed Data
For observed data $\mathcal{D}^{*}$, we approximate the parameter posterior $p(\Theta|\mathcal{D}^{*})$ using Approximate Bayesian Computation (ABC). ABC algorithms follow the basic format of Algorithm 1, which is ABC with rejection sampling. In this section, we discuss the choices of priors $p(\Theta)$, distance function $\nu$, and tolerance $\epsilon$.
Algorithm 1 ABC with rejection sampling
3:for Candidate parameter values $\theta_{i}$ sampled from $p(\Theta)$ do
5:if $\nu(\mathcal{D}^{*},\mathcal{D}_{i})\leq\epsilon$ then
6:  accept $\theta_{i}$
7:end if
8:end for
Choosing priors $p(\Theta)$
For each parameter in each model, we use the uniform prior $p(\theta)=\mathcal{U}_{[\theta_{\text{min}},\theta_{\text{max}}]}(\theta)$, where $\theta_{\text{min}},\theta_{\text{max}}]$ are the parameter ranges in Table 1. Each parameter range is either taken from existing literature or chosen with the aim of exhibiting a wide range of simulation behaviour in each model–see Table 1 for details.
Choosing distance function $\nu$
Following section 9.1, we choose a subset of $n_{\textnormal{s}}=100$ summary statistics for each model from among the spatially-averaged and topological summary statistics. Due to the computational demands of parameter inference using ABC-SMC, it is not computationally feasible to benchmark multiple potential values of $n_{\textnormal{s}}$ or multiple subsets of summary statistics. We therefore advocate a heuristic approach to choosing he number of summary statistics to use in $\nu$. As shown in Figure 9, most of the predictive power of each Random Forest lies in the top $100$ summary statistics. Including additional summary statistics is unlikely to aid parameter inference, as they would only add noise to the ABC distance function $\nu$. Since it is likely that different summary statistics will be informative for different models, the value of $n_{s}$ must be set large enough to ensure that some of the same summary statistics are chosen for all three models. Setting $n_{\textnormal{s}}=100$ results in $\tilde{n}_{\text{s}}=30$ summary statistics selected for all three models, which we deem sufficient to train the Random Forests needed for model selection.
@@
It is possible that the important summary statistics may be poorly scaled. For example, connected components are generally larger than loops in the angiogenesis datasets we consider. Small differences in the size and location of loops in angiogenesis datasets may be stronger predictors of parameter values than the size and locations of connected components, yet this may not be reflected in the distance function $\nu$ if the corresponding topological summary statistics are left unscaled. We therefore construct a scaling function which divides each summary statistic by the maximum absolute value of that summary statistic among the training data for each model. The ABC distance function is then $\nu(\mathcal{D}^{*},\mathcal{D}_{i})=\norm{x^{*}-x_{i}}_{2}$ where $\mathcal{D}^{*}$ is observed data, $\mathcal{D}_{i}$ is simulated data, and $x^{*}$ and $x_{i}$ are the $100$ summary statistics with the scaling applied. When $\mathcal{D}^{*}$ comprises several instances of observed data, we compute the distance $\nu$ for each observed dataset and take their average.
Choosing tolerance $\epsilon$
@@
Using Algorithm 1 with a single value of the tolerance $\epsilon$ may lead to a poor approximate posterior or slow convergence. If $\epsilon$ is too large, too many parameters will be accepted from the prior, and if $\epsilon$ is too small, few parameters will be accepted and many simulations will be needed to approach the true posterior. Instead, we use the ABC-SMC algorithm of smcDelMoral which outputs a series of $n_{\text{pop}}$ intermediate distributions, which correspond to a decreasing sequence of tolerances $\epsilon_{0}>\dots>\epsilon_{n_{\text{pop}}}>0$. Beginning with a population of $N_{\text{pop}}$ parameters sampled from a prior distribution, the initial tolerance $\epsilon_{0}$ is chosen so that a predetermined fraction $\alpha\in(0,1)$ of parameters are accepted. The effective sample size (ESS) of each population is a measure of the independence of parameters within a population. smcDelMoral computes the ESS of each subsequent population and chooses further tolerances such the ESS of each population decreases by $\alpha$. We use $\alpha=0.8$ and generate $n_{\text{pop}}=50$ populations, each containing $N_{\text{pop}}=1,000$ parameters, and take the final population as the approximate ABC-SMC posterior $p(\Theta|\mathcal{D}^{*})$.
@@ 9.3 Step 3: Approximate The Model Posterior
We approximate the model posterior $p(m|\mathcal{D}^{*})$ for observed data $\mathcal{D}^{*}$ by training two more Random Forests, following the method of ABCRF_modelchoice. Using the training data $(X_{i},m_{i})$ obtained from the training data used in step 1 by replacing the parameter $\theta_{i}$ with the model index $m_{i}\in\{\text{AC},\text{SL},\text{PS}\}$, we train a classification Random Forest to learn the relationship between summary statistics $X_{i}$ and model index $m_{i}$.
Step 1 chooses a subset of the spatially-averaged and topological summary statistics which are important for each for each model. Since the importance of each summary statistic to a model is determined by that model’s training data, it is possible that different summary statistics are important for different models. Using summary statistics which are only informative for a single model to approximate the model posterior may bias predictions in favour of that model. We therefore learn the relationship between summary statistics and model index using only those summary statistics which are important for all angiogenesis models.
@@
The trained classification Random Forest provides a prediction $RF(X^{*})$ of the model $m^{*}$ which generated observed data $\mathcal{D}^{*}$ (which we aggregate when $\mathcal{D}^{*}$ contains multiple instances of observed data). The Random Forest also provides an out-of-bag error for each pair $(X_{i},m_{i})$, which is the proportion of those decision trees which did not use this pair in their training bootstrap which predict the incorrect model index $m_{i}$. Following ABCRF_modelchoice, we train another regression Random Forest using the same training data to learn the relationship between summary statistics $X_{i}$ and this mis-classification error rate $p(RF(X_{i})\neq m_{i})$ of the first Random Forest. Predicting the out-of-bag error $p(RF(X^{*})\neq m^{*})$ for the observed data $\mathcal{D}^{*}$ gives an estimate $1-p(RF(X^{*})\neq m^{*})$ for the model posterior $p(m=m^{*}|\mathcal{D}^{*})$.
Figure 9: Left axis: for each parameter in each model, we rank summary statistics by their importance and observe how importance decreases. Right axis: we compare how many summary statistics appear among the top $n_{\text{s}}$ most important features for all models as $n_{\text{s}}$ is increased and plot the expected number of common summary statistics if they were chosen randomly. Choosing $n_{\text{s}}=100$ includes approximately the $25$ most important summary statistics for each parameter. Selecting more summary statistics than this would include those which are of little use in capturing the effect of parameter values on simulated data. Choosing $n_{\text{s}}=100$ also ensures that $\tilde{n}_{\text{s}}=30$ summary statistics appear in the top $n_{\text{s}}$ most important summary statistics for all models. If $100$ summary statistics were chosen uniformly randomly for each model, the expected number that would be common to all three models is approximately 11.